# Hotfixing Large Language Models for Code: How Far Can Parameter-Efficient Fine-Tuning Go?

### v2

@@ 5.3. RQ3: How does hotfixing affect functional correctness and how to mitigate the undesired side effects?
Previous study (Chen et al., 2018) highlights that hotfixes need to remediate the unwanted symptoms of the critical issue and there is less emphasis on correctness. We evaluate how hotfixing LLM4Code introduces undesired side effects on correctness and how to mitigate them. We consider potential side effects on two properties: functional correctness and model’s perplexity, the definitions of which are explained in Section 4.3. Based on the results in RQ1 and RQ2, we focus on the best combination of hotfixing setting: i.e., using the Penalize loss function and LoRA to fine-tuning the models. We additionally consider Penalize+KL to show whether adding KL divergence to the loss function can help mitigate the side effects.
@@
For quantifying the functional correctness, we compute the pass@100 metric on the HumanEval benchmark (Chen et al., 2021). The results are shown in Table 3. We can observe that hotfixing LLM4Code reduces the pass@100 value of both models. When the Penalize loss function is used, the pass@100 value decreases from 16.1 to 14.3 for CodeGen-350M and from 32.9 to 30.4 for CodeGen-2B. But we find that the side effects can be mitigated by adding KL divergence to the loss function. Specifically, using the Penalize+KL loss function can increase the pass@100 value from 14.3 to 14.9 for CodeGen-350M and from 30.4 to 31.1 for CodeGen-2B, compared to the results achieved by Penalize loss function. We also conduct a statistical analysis to evaluate the significance of the difference between the functional correctness of original models and the models after hotfixing using Penalize+KL. The HumanEval benchmark has 164 questions and we compute the pass@100 value for each question for each model. Given two vectors of pass@100 values (obtained by the original models and the hotfixed model using Penalize+KL), we use the Wilcoxon signed-rank test (Wilcoxon, 1945) to compare the two vectors. We find that the $p$-value is larger than 0.05, meaning that although hotfixing LLM4Code can decrease the functional correctness value but the difference is not statistically significant.
@@
We then analyze how the models’ perplexity changes. A smaller perplexity value indicates that the model is more confident predicting each token in the sequence. We follow the settings in Section 4.3 to compute the model perplexity and present the results in Table 3. Similar observations are made on perplexity: hotfixing LLM4Code increases model perplexity but adding KL divergence to the loss function can help mitigate this side effect.
@@
Answers to RQ3: Hotfixing LLM4Code can decrease the functional correctness and increase the model perplexity. Adding KL divergence can mitigate the side effects. When using the Penalize+KL loss function, hotfixing LLM4Code does not significantly affect the functional correctness.
@@
Table 3. How does hotfixing affect the performance of LLM4Code on the HumanEval benchmark and perplexity. We report the pass@100 value (the higher the better) and perplexity (the lower the better) of the original models and the models after hotfixing.
@@
Penalize 14.3 41.37 30.4 3.941
Penalize+KL 14.9 37.19 31.1 3.831
@@ 6. Discussion
We present some discussion in this section, including (1) ethical consideration of the research, (2) potential applications of the hotfix, and (3) threats to validity.
@@ 6.1. Ethical Consideration
The dataset for privacy leakage mitigation is obtained from HuggingFace. [^7] https://huggingface.co/datasets/bigcode/bigcode-pii-dataset-training To access the dataset, one must agree to the terms of usage of this dataset. As explicitly stated by BigCode, the owner of this dataset, the user must agree that you will not share the PII dataset or any modified versions for whatever purpose. To comply with the terms, we do not share the dataset used in our experiments relevant to privacy leakage mitigation. Throughout the research process, we ensure that we handle data and results in an ethical manner. For example, in Listing 2, we obfuscate identifiers and mask the sensitive information so that the contributors of the vulnerable code cannot be disclosed. Our study aims to use hotfix to avoid generating undesired code. However, it is possible that the method is used in the opposite way, i.e., to generate undesired code, which may cause new security and privacy issues. We leave the potential ways to mitigate this risk as future work.
@@ 6.2. Potential Applications
In a nutshell, this paper evaluates the potential of mitigating the undesired behavior of code generation models without heavy retraining. We posit that hotfix has the potential for broad applicability across numerous contexts.
@@
Deployment. A major advantage of the proposed method is its ease of deployment and distribution. Let us consider two typical deployment scenarios: (1) cloud-based deployment (e.g., Copilot) and (2) client-side deployment (e.g., CodeGeeX) In the case of cloud deployment, service providers can simply apply the hotfix by adding the LoRA matrix to certain layers of a network. For client-side deployment, it’s sufficient to distribute only the LoRA matrix, which is relatively small in size (13.9 MB for a CodeGen-2B), to the client. This eliminates the need to download the entire model, which can be significantly larger in size (5.69 GB for a CodeGen-2B). This approach is akin to distributing a software patch as opposed to the entire software package, greatly simplifying the deployment process of the proposed method.
@@ 6.3.1. Threats to Internal Validity
Internal validity refers to the extent to which a study’s results are not biased or derived wrongly. When evaluating the SStuBs-fixing task, we follow the setting used by Jesse et al. (Jesse et al., 2023) that evaluates whether a model can generate fixed code by string matching. However, it is possible that the model generates code that is semantically equivalent to the fixed code but not exactly the same. On the privacy leakage mitigation task, we need to detect whether privacy information appears in the generated code and the existing privacy detection tools are not perfect. To mitigate this threat, we use a state-of-the-art privacy detection tool (Bigcode/Starpii) and only focus on emails, which are the type of PII on which the tool performs best to identify (with a precision of 97.73%). Another threat is that the deep learning models are nondeterministic, which means that the same model may generate different outputs on the same prompt. To mitigate this threat, we run each model 10 times on each prompt in the SStuBs-fixing task and 20,000 times in the privacy leakage mitigation task.
@@ 6.3.2. Threats to External Validity
External validity refers to how the results of this study generalize to other settings. For example, the conclusion may not hold for other models. Our experiments investigate the CodeGen model family, which is widely used and is based on the GPT-2 architecture. We leave the evaluation of encoder-decoder models (e.g., CodeT5) as future work. To mitigate the threats that the conclusion may vary for different sizes of models, we choose two models of different sizes. We mitigate the undesired behavior: LLM4Code can complete many known buggy code, aiming at make LLM4Code not complete buggy code but complete the corresponding fixed code. To evaluate the generalizability of the proposed method, we conduct a case study on privacy leakage mitigation. Additionally, to evaluate the functional correctness of LLM4Code after hotfixing, we use the HumanEval benchmark (Chen et al., 2021), which is a widely used benchmark for evaluating code generation models. However, this benchmark only contains Python programming problems, which leads to a threat that the conclusion may not hold for other programming languages. We leave the evaluation of other languages as future work.
@@ 7. Related Work
This is an extended replication study of the work by He and Vechev, which use contrastive prefix-tuning (Qian et al., 2022) to control the security-level of code generators. We extend their work on a new hotfixing task (i.e., reduce the number of generated buggy code and increase the number of fixed code) with more parameter-effect fine-tuning methods. We discuss the related works of other undesired behaviors in LLM4Code and efficient-tuning methods for models of code as follows.
@@ 7.1. LLM4Code and Behaviors
In addition to their numerous applications, researchers have identified key limitations and undesired behaviors in LLM4Code (Lo, 2023; Yang et al., 2024a). For example, Yang et al. (Yang et al., 2024b) and Al-Kaswan et al. (Al-Kaswan et al., 2023) find that code models can memorize training data, which can lead to privacy leakage. Similarly, researchers find the LLM4Code can generate software credentials like API keys and passwords (Huang et al., 2023b; Niu et al., 2023). Jesse et al. (Jesse et al., 2023) find that code models tend to generate buggy code. A series of empirical studies show that code generators can produce biased (Huang et al., 2024; Liu et al., 2023) and vulnerable code (Pearce et al., 2022; Sandoval et al., 2022). To the best of our knowledge, we are the first to hotfix to mitigate the above two types of undesired behaviors: generating buggy code and leaking privacy information. There are other undesired behaviors in LLM4Code. Huang et al. (Huang et al., 2024) find that LLM4Code exhibit social bias in code generation. Besides, a series of studies have shown that LLM4Code are not robust (Yang et al., 2022), vulnerable to data poisoning (Ramakrishnan and Albarghouthi, 2022; Yang et al., [n. d.]; Wan et al., 2022; Li et al., 2022), and lack explainability (Liu et al., 2023), which are not desired as well. We leave the mitigation of these undesired behaviors as future work.
@@ 7.2. Efficient-tuning for Models of Code
Weyssow et al. (Weyssow et al., 2024) use 5 parameter-efficient methods to fine-tune models for code generation tasks. We reuse their replication package to implement the first hotixing for mitigating undesired behavior. Prefix-tuning falls into the category of efficient-tuning methods for pre-trained models, which aim to utilize computationally-friendly techniques to efficiently adapt a model to a new dataset or task. Typical methods include in-context learning, prefix-tuning, adapter-tuning. Here we briefly introduce the related work about applying these methods to models of code. In-context learning (ICL) presents a small number of examples in the prompt to guide language models to conduct specific tasks. Gao et al. (Gao et al., 2023) evaluate ICL on code summarization, bug fixing, and program synthesis. They show that strategically designed context can lead to substantial improvements over widely-used demonstration construction methods. Huang et al. (Huang et al., 2023a) show that ICL can mitigate the social bias of code generated by language models. Wang et al. (Wang et al., 2022) apply prompt-tuning on three code intelligence tasks and show that prompt-tuning consistently outperforms fine-tuning in three investigated tasks. Wang et al. (Wang et al., 2023) evaluate adapter tuning for code search and summarization, showing that adapter tuning significantly outperforms full-model fine-tuning and effectively overcomes catastrophic forgetting. Choi and Lee (Choi and Lee, 2023) propose CodePrompt, using task-agnostic prefix tuning for program and language generation.
@@ 8. Conclusion and Future Work
In this study, we explore to what extent parameter-efficient fine-tuning (PEFT) techniques can be used to hotfix LLM4Code. We extend the work by He and Vechev (Qian et al., 2022) by evaluating four different PEFT techniques on the task of reducing the number of generated buggy code and increasing the number of fixed code. We conduct our experiments on the models in the CodeGen family. The empirical results demonstrate the effectiveness of hotfixing: the best combination leads to a significant increase in the generation of fixed code (up to 108.42%) and a notable reduction in the production of buggy code (up to 50.47%). We also conduct statistical analysis to confirm that hotfixing does not adversely impact the functional performance of the models on the HumanEval benchmark. We discuss the potential applications of the proposed method in various tasks, as well as its advantage in efficient deployment.
In the future, we plan to extend hotfixing to more undesired behaviors and LLM4Code of various sizes and architectures.

### v4

@@ 4. Results
Previous studies (Chen et al., 2018) highlight that hotfixes need to remediate the unwanted symptoms of critical issues, with less emphasis on correctness. We evaluate how hotfixing LLM4Code introduces undesired side effects on correctness and how to mitigate them. We consider potential side effects on two properties: functional correctness and model perplexity, the definitions of which are explained in Section 3.6. Based on the results in RQ1 and RQ2, we focus on the best combination of hotfixing settings: i.e., using the Duel loss function and LoRA to fine-tune the models. We additionally consider Duel+KL to show whether adding KL divergence to the loss function can help mitigate the side effects.
@@
For quantifying the functional correctness, we compute the pass@100 metric on the HumanEval benchmark (Chen et al., 2021). The results are shown in Table 4. We can observe that hotfixing LLM4Code reduces the pass@100 value of both models. When the Duel loss function is used, the pass@100 value decreases from 16.1 to 14.3 for CodeGen-350M and from 32.9 to 30.4 for CodeGen-2B. However, we find that the side effects can be mitigated by adding KL divergence to the loss function. Specifically, using the Duel+KL loss function can increase the pass@100 value from 14.3 to 14.9 for CodeGen-350M and from 30.4 to 31.1 for CodeGen-2B, compared to the results achieved by the Duel loss function. We also conduct a statistical analysis to evaluate the significance of the difference between the functional correctness of the original models and the models after hotfixing using Duel+KL. The HumanEval benchmark has 164 questions, and we compute the pass@100 value for each question for each model. Given two vectors of pass@100 values (obtained by the original models and the hotfixed model using Duel+KL), we use the Wilcoxon signed-rank test (Wilcoxon, 1945) to compare the two vectors. We find that the $p$-value is larger than 0.05, meaning that although hotfixing LLM4Code can decrease the functional correctness value, the difference is not statistically significant.
@@
We then analyze how the models’ perplexity changes. A smaller perplexity value indicates that the model is more confident in predicting each token in the sequence. We follow the settings in Section 3.6 to compute the model perplexity and present the results in Table 4. Similar observations are made on perplexity: hotfixing LLM4Code increases model perplexity, but adding KL divergence to the loss function can help mitigate this side effect.
@@
Answers to RQ3: Hotfixing LLM4Code can decrease functional correctness and increase model perplexity. Adding KL divergence can mitigate the side effects. When using the Duel+KL loss function, hotfixing LLM4Code does not significantly affect functional correctness.
@@
Table 4. How does offline hotfixing affect the performance of LLM4Code on the HumanEval benchmark and perplexity? We report the pass@100 value (the higher the better) and perplexity (the lower the better) of the original models and the models after hotfixing.
@@
Duel 14.3 41.37 30.4 3.941
Duel+KL 14.9 37.19 31.1 3.831
5. A Case Study on Privacy Mitigation
To demonstrate the generalizability of hotfixing, we conduct a case study to mitigate another type of undesired behavior recently exposed by researchers: sensitive information leakage. Researchers (Feng et al., 2022; Basak et al., 2023) have shown that public repositories contain a large amount of sensitive information, including personally identifiable information (PII), like email addresses and human names. Yang et al. (Yang et al., 2024c) find that code generators memorize such information from the training data and generate code that exposes privacy. An example is shown in Listing 2, where a code model memorizes a hardcoded public IP address, username, and password from its training data.
1netowrk_config = {
2 ’device_type’: <masked value>,
3 ’ip’: <masked value>,
4 ’username’: <masked value>,
5 ’password’: <masked value>,
6}
7network_con = ConnectHandler(**netowrk_config)
8print network_con.find_prompt()
Listing 2: An example of generated code that contains sensitive information, e.g., public IP address, username, etc. We have substituted the identifiers and strings with placeholders to preserve privacy.
We leverage the method used in Yang et al.’s study (Yang et al., 2024c) to evaluate to what extent CodeGen-350M and CodeGen-2B memorize and produce sensitive information. We feed the ‘start-of-sequence’ token into the model and let the model produce outputs in an auto-regressive manner. By default, each model generates 20,000 outputs with 512 tokens each. We run the state-of-the-art privacy detection tool starpii (Li et al., 2023) on the generated outputs. The starpii tool conducts a named entity recognition (NER) task, classifying each character in an input. The model starpii has a precision of 52.20% and 70.94% in detecting usernames and passwords, respectively. In contrast, the precision of detecting emails is 97.73%. To better evaluate how the proposed method can help models mitigate privacy leakage, we choose to focus on emails. We count the occurrence of detected emails in the generated outputs and compute the email leakage ratio, i.e., how many emails are leaked per generation on average. We find that the email leakage ratio is 0.0859 for CodeGen-350M and 0.0833 for CodeGen-2B; results are shown in Table 5.
We employ offline hotfixing for its high effectiveness. To hotfix LLM4Code, we need code snippets and the annotated privacy information in the code snippets. We use the dataset hosted on HuggingFace: bigcode/bigcode-pii-dataset, [^6] https://huggingface.co/datasets/bigcode/bigcode-pii-dataset which is released in (Li et al., 2023). This is an annotated dataset for Personally Identifiable Information (PII) in code. The annotation process involved 1,399 crowd-workers from 35 countries with Toloka, [^7] https://toloka.ai a platform that supports crowdsourcing data annotation. This dataset consists of 12,099 samples of around 50 lines of code in 31 programming languages. There are multiple types of PII annotated in the dataset, including emails, IP addresses, usernames, passwords, keys, etc. However, this paper focuses on emails for accurate evaluation considerations. After filtering out the samples that do not contain emails, we obtain 640 examples for training. Given a code and its annotation, showing which part of the code is PII, we replace the PII part with placeholders. Specifically, we replace emails with ‘<EMAIL>’, producing a new version of the code. In this way, we obtain the diff between the original code and the code with placeholders, which is used to train the hotfix.
Table 5. How hotfixing affect email leakage ratios.
Category Model Original Hotfix Reduction
No. of Emails CodeGen-350M 0.0859 0.0006 99.30%
CodeGen-2B 0.0833 0.0017 97.96%
Based on the observations made in previous RQs, we use LoRA as the fine-tuning method and Duel+KL as the loss function. From Table 5, we can observe that hotfixing reduces the exposure of emails by 99.30% and 97.96% on CodeGen models with 350M and 2B parameters, respectively. It suggests that the hotfix provides better privacy mitigation on smaller models.
We also explore whether the results can generalize when a model generates outputs of different lengths. We let the CodeGen-350M model generate 3 more sets of outputs, each containing 20,000 outputs. The three maximal tokens of outputs in the 3 sets are set as 256, 768, and 1,024, following the settings in Yang et al.’s experiment (yang2023memorization). We find that the privacy reduction rates are 99.18%, 99.30%, 98.42%, and 98.58%, showing the potential of hotfixing in mitigating privacy leakage. Note that hotfixing CodeGen-350M only takes 5 minutes, as the training data is small, having 640 examples with annotated emails.
Findings from the case study: Hotfixing can reduce the number of generated emails by up to 99.30%. The results can be generalized when the model generates outputs of different lengths.
@@ 6. Discussion
We present some discussions in this section, including (1) ethical considerations of the research, (2) potential applications of the hotfix, and (3) threats to validity.
@@ 6.1. Ethical Consideration
The dataset for privacy leakage mitigation is obtained from HuggingFace. [^8] https://huggingface.co/datasets/bigcode/bigcode-pii-dataset-training To access the dataset, one must agree to the terms of usage of this dataset. As explicitly stated by BigCode, the owner of this dataset, the user must agree that you will not share the PII dataset or any modified versions for whatever purpose. To comply with the terms, we do not share the dataset used in our experiments relevant to privacy leakage mitigation. Throughout the research process, we ensure that we handle data and results in an ethical manner. For example, in Listing 2, we obfuscate identifiers and mask the sensitive information so that the contributors of the vulnerable code cannot be disclosed. Our study aims to use hotfixes to avoid generating undesired code. However, it is possible that the method is used in the opposite way, i.e., to expose personal information, which may cause new privacy concerns. We leave the potential ways to mitigate this risk as future work.
@@ 6.2. Potential Applications
In a nutshell, this paper evaluates the potential of mitigating the undesired behavior of code generation models without extensive retraining. We posit that hotfixes have the potential for broad applicability across numerous contexts.
@@
Deployment. A major advantage of the proposed method is its ease of deployment and distribution. Let us consider two typical deployment scenarios: (1) cloud-based deployment (e.g., Copilot) and (2) client-side deployment (e.g., CodeGeeX). In the case of cloud deployment, service providers can simply apply the hotfix by adding the LoRA matrix to certain layers of a network. For client-side deployment, it’s sufficient to distribute only the LoRA matrix, which is relatively small in size (13.9 MB for a CodeGen-2B), to the client. This eliminates the need to download the entire model, which can be significantly larger in size (5.69 GB for a CodeGen-2B). This approach is akin to distributing a software patch as opposed to the entire software package, greatly simplifying the deployment process of the proposed method.
@@ 6.3.1. Threats to Internal Validity
Internal validity refers to the extent to which a study’s results are not biased or derived wrongly. When evaluating the SStuBs-fixing task, we follow the setting used by Jesse et al. (Jesse et al., 2023) that evaluates whether a model can generate fixed code by string matching. However, it is possible that the model generates code that is semantically equivalent to the fixed code but not exactly the same. On the privacy leakage mitigation task, we need to detect whether privacy information appears in the generated code, and the existing privacy detection tools are not perfect. To mitigate this threat, we use a state-of-the-art privacy detection tool (Bigcode/Starpii) and only focus on emails, which are the type of PII on which the tool performs best to identify (with a precision of 97.73%). Another threat is that the deep learning models are nondeterministic, which means that the same model may generate different outputs on the same prompt. To mitigate this threat, we run each model 10 times on each prompt in the SStuBs-fixing task and 20,000 times in the privacy leakage mitigation task.
@@ 6.3.2. Threats to External Validity
External validity refers to how the results of this study generalize to other settings. For example, the conclusion may not hold for other models. Our experiments investigate the CodeGen model family, which is widely used and is based on the GPT-2 architecture. We leave the evaluation of encoder-decoder models (e.g., CodeT5) as future work. To mitigate the threats that the conclusion may vary for different sizes of models, we choose two models of different sizes. We mitigate the undesired behavior: LLM4Code can complete many known buggy codes, aiming to make LLM4Code not complete buggy code but complete the corresponding fixed code. To evaluate the generalizability of the proposed method, we conduct a case study on privacy leakage mitigation. Additionally, to evaluate the functional correctness of LLM4Code after hotfixing, we use the HumanEval benchmark, which is a widely used benchmark for evaluating code generation models. However, this benchmark only contains Python programming problems, which leads to a threat that the conclusion may not hold for other programming languages. We leave the evaluation of other languages as future work.
@@ 7. Related Work
In this section, we present an overview of studies relevant to this paper. We use the keyword combination “hotfix” + “large language model” to conduct a literature search on IEEE Xplore and ACM Digital Library, which returns no matching papers. This indicates that hotfixing has not been proposed for LLMs. We select the strategy used by Jesse et al. (Jesse et al., 2023) and the work by He and Vechev (He and Vechev, 2023) as representative for online and offline hotfixing. We also extend them for more detailed analysis on LLM4Code hotfixing. We specifically focus on two hotfixing tasks: bug-fixing and privacy leakage mitigation. We discuss the related works on other undesired behaviors in LLM4Code and efficient tuning methods for models of code as follows.
@@ 7.1. LLM4Code and Behaviors
In addition to their numerous applications, researchers have identified key limitations and undesired behaviors in LLM4Code (Lo, 2023; Yang et al., 2024a). For example, Yang et al. (Yang et al., 2024c) and Al-Kaswan et al. (Al-Kaswan et al., 2024) find that code models can memorize training data, which can lead to privacy leakage. Similarly, researchers find that LLM4Code can generate software credentials like API keys and passwords (Huang et al., 2023b; Niu et al., 2023). Jesse et al. (Jesse et al., 2023) find that code models tend to generate buggy code. A series of empirical studies show that code generators can produce biased (Huang et al., 2024; Liu et al., 2023) and vulnerable code (Pearce et al., 2022; Sandoval et al., 2022). To the best of our knowledge, we are the first to offer a hotfix to mitigate the above two types of undesired behaviors: generating buggy code and leaking privacy information. There are other undesired behaviors in LLM4Code. Huang et al. (Huang et al., 2024) find that LLM4Code exhibit social bias in code generation. Besides, a series of studies have shown that LLM4Code are not robust (Yang et al., 2022), are vulnerable to data poisoning (Ramakrishnan and Albarghouthi, 2022; Yang et al., 2024b; Wan et al., 2022; Li et al., 2024a), and lack explainability (Liu et al., 2023), which are not desirable as well. We leave the mitigation of these undesired behaviors as future work.
@@ 7.2. Efficient-tuning for Models of Code
Weyssow et al. (Weyssow et al., 2024) use 5 parameter-efficient methods to fine-tune models for code generation tasks. We reuse their replication package to implement the first hotfixing for mitigating undesired behavior. Prefix-tuning falls into the category of efficient-tuning methods for pre-trained models, which aim to utilize computationally-friendly techniques to efficiently adapt a model to a new dataset or task. Typical methods include in-context learning, prefix-tuning, and adapter-tuning. Here we briefly introduce the related work about applying these methods to models of code. In-context learning (ICL) presents a small number of examples in the prompt to guide language models to conduct specific tasks. Gao et al. (Gao et al., 2023) evaluate ICL on code summarization, bug fixing, and program synthesis. They show that strategically designed contexts can lead to substantial improvements over widely used demonstration construction methods. Huang et al. (Huang et al., 2023a) show that ICL can mitigate the social bias of code generated by language models. Wang et al. (Wang et al., 2022) apply prompt-tuning on three code intelligence tasks and show that prompt-tuning consistently outperforms fine-tuning in the three investigated tasks. Wang et al. (Wang et al., 2023) evaluate adapter tuning for code search and summarization, showing that adapter tuning significantly outperforms full-model fine-tuning and effectively overcomes catastrophic forgetting. Choi and Lee (Choi and Lee, 2023) propose CodePrompt, using task-agnostic prefix tuning for program and language generation.
@@ 8. Conclusion and Future Work
In this study, we proposed the novel concept of hotfixing LLM4Code—a model maintenance activity aimed at implementing time-critical improvements that are temporary, small in size, fast to obtain, and target specific issues in LLM4Code. We investigated both online and offline hotfixing strategies to adapt LLM4Code to critical changes in the software environment, such as generating fixed code instead of buggy code and producing privacy-preserving completions. Our experiments extended existing strategies and evaluated parameter-efficient fine-tuning (PEFT) techniques, including LoRA (Hu et al., 2021), on the CodeGen family of models. The empirical results demonstrate the effectiveness of hotfixing: using LoRA, we increased the generation of fixed code by up to 108.42% and reduced the generation of buggy code by up to 50.47%. We also achieved a reduction of up to 99.30% in email leakage, enhancing the privacy-preserving capabilities of the models. Importantly, our statistical analysis confirmed that hotfixing does not adversely impact the functional performance of the models on benchmarks like HumanEval.
We believe that hotfixing has broad potential applications, such as updating models for deprecated APIs and customizing code generation for specific development practices. Future work includes extending hotfixing techniques to address more undesired behaviors and applying them to a wider range of LLM4Code models of various sizes and architectures.
For peer review purposes, the replication package is provided at https://anonymous.4open.science/r/Hotfixing-LLM4Code. We will make the replication package publicly available.